OpenStudy (anonymous):

Find length of perimeter of curve x=2 cos^3 t, y= 2 sin^3 t

OpenStudy (amistre64):

x=2 cos^3(t), y= 2 sin^3(t) x'= -6 cos^2(t) sin(t), y' = 6 sin^2(t) cos(t) \(\int_{a}^{b}\sqrt{(-6 cos^2(t) sin(t))^2+(6 sin^2(t) cos(t))^2}\)dt maybe? Since we only have 2 components; we dont need the f(x) part right?

OpenStudy (amistre64):

its a star :)

OpenStudy (anonymous):

Yup, bit of symmetry there....

OpenStudy (amistre64):

so i guess we can get away with 4* int[0,2]

OpenStudy (anonymous):

pi/2

OpenStudy (amistre64):

well, that too

OpenStudy (anonymous):

Works out exactly if anyone is looking to practice.....

OpenStudy (anonymous):

I have a question, in "Calc 111" they teach line/surface integrals with vector formulations as well?

OpenStudy (amistre64):

If its right; 24cos(t)sin(t), [0,pi/2] ??

OpenStudy (amistre64):

that wouldnt work out well at all .... 0 - 0 = 0 lol

OpenStudy (amistre64):

might be 12 in the end

OpenStudy (anonymous):

Yes, 12.......(2 sint cost -> sin 2t...)

OpenStudy (amistre64):

\[4\int_{a}^{b}\sqrt{(-6 cos^2(t) sin(t))^2+(6 sin^2(t) cos(t))^2}dt\] \[4\int_{a}^{b}\sqrt{(36 cos^2(t)cos^2(t) sin^2(t) +36 sin^2(t)sin^2(t) cos^2(t)}dt\] \[4\int_{a}^{b}\sqrt{36 cos^2(t)sin^2(t)\ (cos^2(t)+sin^2(t))}dt\] \[4\int_{a}^{b}(6 sin(t))cos(t)dt\] \[4*3 sin^2(t)dt;\ [0,pi/2]\]

OpenStudy (amistre64):

12(1) - 12(0) = 12

OpenStudy (amistre64):

the first time i simplified the sqrt and forgot to actually integrate lol

OpenStudy (anonymous):

That will do as well..

OpenStudy (lalaly):

ESTUDIER COME BACK :(